OpenStudy (anonymous):

2.Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770?

OpenStudy (shane_b):

It's another simultaneous linear equation problem. First, set up the two equations. Do you know how to do that?

OpenStudy (anonymous):

i dont want the answer, i just want to know how to set up the problem.

OpenStudy (shane_b):

I realize that...you start by creating the two equations. \[x+y=150\]\[45x+28.5y=4770\]Do you see how that was obtained?

OpenStudy (anonymous):

try to see if this is correct set up: 45x+28.5y=4770 and x+y=150 You can rearrange equation two to get: x=150-y You can substitute this into equation one to get: 45(150-y)+28.5y=4770 Expand the brackets and simplify: 6750-45y+28.5y=4770 6750-16.5y=4770 -16.5y=-1980 16.5y=1980 y=120 This means there were 120 students on the trip. x+y=150 x+120=150 x=30

OpenStudy (anonymous):

Yes shane, thank you. and thanks rasmus.

OpenStudy (shane_b):

I guess you got your answer anyway...if you don't understand any of the steps please ask.

OpenStudy (anonymous):

no problem mik.

OpenStudy (anonymous):

alright, and thanks again (: